# Implicit Bias in Matrix Factorization and its Explicit Realization in a New Architecture

### v1

@@ B.1 Design Variants of the UDV Architecture
We only report results from UDV, as defined in equation (7), in the main text. However, we provide results for the other variants in the Appendices for completeness and to illustrate the impact of different constraint settings on model performance.
Figure SM11 and Figure SM12 show that UDV consistently finds a low-rank solution. Generally, UDV exhibits the most pronounced decaying pattern in singular values. Additionally, we extended our experiments to include full-batch training on the HPART and NYCTTD datasets. Although full-batch training converges more slowly than stochastic (mini-batch) methods, it exhibits a similar singular value decay pattern, confirming our earlier observations.
B.2 Comparison of Model Performance with the MBGDM Algorithm
In Section 4.1, we presented the performance of each model (UDV, UV, and UV-ReLU) in terms of generalization power (measured by validation loss or validation accuracy) along with the corresponding singular value spectrum, as shown in Table 1 and Figure 3. The results were obtained by selecting the optimal algorithm and learning rate pair for each model. Here, we conduct the same experiment, but focus exclusively on the MBGDM algorithm. The results, presented in Table SM1 and Figure SM10 are similar to the ones in Section 4.1, showing similar trends, highlighting consistency across methods.
Table SM1: Model performance using MBGDM. M, E, and R represent the transferred models MaxVit-T, EfficientNet-B0, and RegNetX-32GF, respectively.
Tasks Regression (Test Loss) Classification (Test Accuracy)
Dataset HPART NYCTTD MNIST
UDV
$1.345\times 10^{-3}$
MBGDM: $10^{-1}$
$5.248\times 10^{-6}$
MBGDM: $10^{-1}$
M: 99.67%
MBGDM: $10^{-2}$
E: 99.63%
MBGDM: $10^{-1}$
R: 99.74%
MBGDM: $10^{-2}$
UV
$1.337\times 10^{-3}$
MBGDM: $10^{-2}$
$5.259\times 10^{-6}$
MBGDM: $10^{-1}$
M: 99.69%
MBGDM: $10^{-2}$
E: 99.59%
MBGDM: $10^{-1}$
R: 99.56%
MBGDM: $10^{-1}$
UV-ReLU
$1.244\times 10^{-3}$
MBGDM: $10^{-1}$
$5.264\times 10^{-6}$
MBGDM: $10^{-1}$
M: 99.63%
MBGDM: $10^{-2}$
E: 99.68%
MBGDM: $10^{-1}$
R: 99.66%
MBGDM: $10^{-1}$
Figure SM10: Singular value spectrum corresponding to Table SM1.
@@ B.3 Additional Details on the Pruning Experiment
The SVD-based pruning experiments yield conclusions consistent with the singular value decay pattern. UDV typically exhibits the fastest decay, leading to more compact models while maintaining performance, as shown in Figure SM13.
Figure SM11: Comparison of singular value pattern among all UDV variants, UV-ReLU and UV on the regression tasks.
Figure SM12: Comparison of singular value pattern among all UDV variants, UV-ReLU and UV on the MNIST dataset.
Figure SM13: The performance of SVD-based pruning. The index in x-axis represents the number of neurons in the diagonal layer after pruning. For the HPART and NYCTTD datasets, the validation loss change indicates how much worse the pruned model performs compared to the baseline (the model before pruning), expressed as a percentage ($\frac{loss_{pruned}-loss_{baseline}}{loss_{baseline}}\times 100\%$). Note that the pruned model may outperform the baseline, but negative values cannot be displayed on a logarithmic scale. The red dashed line denotes the 0.1% threshold, indicating negligible performance sacrifice. For the MNIST dataset, the results show the validation accuracy after pruning the model.
We observed that the learning rate can have some impact on the singular value decay pattern. In particular, a very large learning rate may cause oscillations in both training and validation loss, yet may result in a rapid decay of the spectrum. Conversely, a small learning rate may lead to a less pronounced spectral decay but still can yield a comparable validation loss to that of the optimal learning rates. For example, the difference in validation losses between the Adam optimizer with learning rates of 1e-3 and 1e-4 was negligible, yet the singular value spectra differed (see Figure SM14). This discrepancy also affects the performance of the pruning experiments (see Figure SM15).
Figure SM14: Differences in singular value patterns across varying learning rates.
Figure SM15: Different learning rates lead different pruning performance. the validation loss change indicates how much worse the pruned model performs compared to the baseline (the model before pruning), expressed as a percentage ($\frac{loss_{pruned}-loss_{baseline}}{loss_{baseline}}\times 100\%$). Note that the pruned model may outperform the baseline, but negative values cannot be displayed on a logarithmic scale. The red dashed line denotes the 0.1% threshold, indicating negligible performance sacrifice.
When the learning rate is close to the optimal value, a ‘stair-step’ pattern (see Figure SM16) can be observed on the loss or accuracy curve in the classification task. It could be attributed to the model continuously searching for a low-rank solution, increasing the rank gradually.
Figure SM16: The ‘stair-step’ accuracy curve in classification (RegNetX-32GF with Adam). The sub-figures in the second row correspond to the respective singular value spectra, indicating that UDV consistently seeks low-rank solutions regardless of the learning rate.
@@ B.4 Incorporating ReLU into UDV
We conducted a preliminary experiment to verify the feasibility of incorporating ReLU into the UDV, leaving a comprehensive study to future work. We derived the optimizers and learning rates from Tables SM2, SM3, SM4 and SM5, but only used the HPART dataset for regression and the MNIST for classification.
Figure SM17 and Figure SM18 showed that UDV-ReLU and UDV-ReLU-s exhibit a similar (or even more pronounced) decaying pattern in singular values as observed in the proposed UDV. The ‘stair-step’ accuracy curve was also observed in these models.
Figure SM17: ReLU in UDV: Singular value pattern on HPART dataset.
Figure SM18: ReLU in UDV: Validation accuracy (the first row) and the singular value pattern (the second row) on MNIST dataset. Faster convergence can be achieved by fine-tuning the learning rate.
@@ B.5 Comparison of UDV and Three-Layer Fully Connected Networks
The results, also included in Tables SM2, SM3, SM4 and SM5, are shown in Figure SM19 and Figure SM20. These findings demonstrate that the singular value decay in UDV unconstrained, UFV, and UV is significantly slower than in the proposed UDV framework. This underscores the critical role of the diagonal layer and constraints in facilitating low-rank solution identification within the proposed structure.
Figure SM19: Comparison of singular value spectrum among UDV, UDV (unconstrained),UFV and UV on the regression tasks.
Figure SM20: Comparison of singular value spectrum among UDV, UDV (unconstrained), UFV and UV on the classification datasets (MNIST).
@@ B.6 Comparing the Effects of UDV and Weight Decay Regularization
Increasing $\gamma$ led to a faster decay in the singular value spectrum but at the cost of higher training and validation losses. Figure SM21 illustrates an example where we selected the largest $\gamma$ values for which the validation loss remained within 10% of the UDV baseline. The resulting singular value decay in the UV and UFV models was less pronounced than in the UDV model. Further increasing $\gamma$ to match or exceed the spectral decay of the UDV model resulted in significantly worse validation loss compared to the UDV network. The complete results are presented in Figures SM22 and SM23.
@@
Figure SM21: Comparison of singular value spectrum among the proposed UDV, UFV and UV (with weight decay) on the HPART dataset. In addition to the UDV baseline, the UFV and UV models include the regularization parameter $\gamma$ and the relative change in validation performance. The relative change is defined as ($\frac{loss_{model}-loss_{baseline}}{loss_{baseline}}\times 100\%$), representing how much worse the model performs compared to the UDV baseline.
@@
Figure SM22: Extension of Figure SM21. Comparison of singular value spectrum between the proposed UDV and UFV (with weight decay) on the HPART dataset. Regularization parameter $\gamma=0$ indicates that no weight decay is applied.
@@
Figure SM23: Extension of Figure SM21. Comparison of singular value spectrum between the proposed UDV and UV (with weight decay) on the HPART dataset. Regularization parameter $\gamma=0$ indicates that no weight decay is applied.
@@
Table SM2: Experiments using Adam optimizer. Not applicable or results with obvious oscillations or divergence are denoted as ‘–’.
@@
(MaxVit-T [M] |||| EfficientNet-B0 [E] |||| RegNetX-32GF [R])
(×$\times 100\%$)
@@
Table SM3: Experiments using NAdam optimizer. Not applicable or results with obvious oscillations or divergence are denoted as ‘–’.
@@
(MaxVit-T [M] |||| EfficientNet-B0 [E] |||| RegNetX-32GF [R])
(×$\times 100\%$)
@@
Table SM4: Experiments using MBGD optimizer. Not applicable or results with obvious oscillations or divergence are denoted as ‘–’.
@@
(MaxVit-T [M] |||| EfficientNet-B0 [E] |||| RegNetX-32GF [R])
(×$\times 100\%$)
@@
Table SM5: Experiments using MBGDM optimizer. Not applicable or results with obvious oscillations or divergence are denoted as ‘–’.
@@
(MaxVit-T [M] |||| EfficientNet-B0 [E] |||| RegNetX-32GF [R])
(×$\times 100\%$)

### v2

@@ B.2 Design Variants of the UDV Architecture
Early image classification experiments were performed on the normalized MNIST dataset (LeCun et al., 2010), which allowed rapid iteration and efficient validation of the proposed method. The replacement of the classifier follows the procedure described in Section 3.1. The number of hidden neurons in the diagonal layer was defined as $m=\texttt{floor}(\tfrac{2}{3}d)$,and this results in a UDV network structure ($d$-$m$-$c$) of 512-341-10 for MaxViT-T, 1280-853-10 for EfficientNet-B0, and 2520-1680-10 for RegNetX-32GF.
Four optimizers, Adam, NAdam, MBGD, and MBGDM, were applied to all tasks. Specifically, for regression tasks, we tested learning rates of ($10^{-4}$, $10^{-3}$, $10^{-2}$, $0^{-1}, 11$, 22 33). Larger learning rates of ($1$, 22, 33) were often excluded for the UV model due to divergence. For classification tasks, we tested learning rates of ($10^{-6}$, $10^{-5}$, $10^{-4}$, $10^{-3}$, $10^{-2}$, $0^{-1}, 11$) with Adam and NAdam, and ($10^{-3}$, $10^{-2}$, $0^{-1}, 11$, $2$, 33, 55) with MBGD and MBGDM.
Model performances are summarized in Tables SM3, SM4, SM5 and SM6, and Figures SM17 and SM18 illustrate that UDV consistently converges to a low-rank solution. Generally, UDV exhibits the most pronounced decaying pattern in singular values. Additionally, we extended our experiments to include full-batch training on the HPART and NYCTTD datasets. Although full-batch training converges more slowly than stochastic (mini-batch) methods, it exhibits a similar singular value decay pattern, confirming our earlier observations.
Figure SM17: Comparison of singular value pattern among all UDV variants, UV-ReLU and UV on the regression tasks.
Figure SM18: Comparison of singular value pattern among all UDV variants, UV-ReLU and UV on the MNIST dataset.
@@ B.3 Additional Details on the Pruning Experiment
The SVD-based pruning experiments yield conclusions consistent with the singular value decay pattern. UDV typically exhibits the fastest decay, leading to more compact models while maintaining performance, as shown in Figure SM19.
Figure SM19: The performance of SVD-based pruning. The index in x-axis represents the number of neurons in the diagonal layer after pruning. For the HPART and NYCTTD datasets, the test loss change indicates how much worse the pruned model performs compared to the baseline (the model before pruning), expressed as a percentage ($\frac{loss_{pruned}-loss_{baseline}}{loss_{baseline}}\times 100\%$). Note that the pruned model may outperform the baseline, but negative values cannot be displayed on a logarithmic scale. The red dashed line denotes the 0.1% threshold, indicating negligible performance sacrifice. For the MNIST dataset, the results show the test accuracy after pruning the model.
We observed that the learning rate can have some impact on the singular value decay pattern. In particular, a very large learning rate may cause oscillations in both training and test loss, yet may result in a rapid decay of the spectrum. Conversely, a small learning rate may lead to a less pronounced spectral decay but still can yield a comparable test loss to that of the optimal learning rates. For example, the difference in test losses between the Adam optimizer with learning rates of $10^{-3}$ and $10^{-4}$ was negligible, yet the singular value spectra differed (see Figure SM20). This discrepancy also affects the performance of the pruning experiments (see Figure SM21).
Figure SM20: Differences in singular value patterns across varying learning rates.
Figure SM21: Different learning rates lead different pruning performance. the test loss change indicates how much worse the pruned model performs compared to the baseline (the model before pruning), expressed as a percentage ($\frac{loss_{pruned}-loss_{baseline}}{loss_{baseline}}\times 100\%$). Note that the pruned model may outperform the baseline, but negative values cannot be displayed on a logarithmic scale. The red dashed line denotes the 0.1% threshold, indicating negligible performance sacrifice.
When the learning rate is close to the optimal value, a ‘stair-step’ pattern (see Figure SM22) can be observed on the loss or accuracy curve in the classification task. It could be attributed to the model continuously searching for a low-rank solution, increasing the rank gradually.
Figure SM22: The ‘stair-step’ accuracy curve in classification (RegNetX-32GF with Adam). The sub-figures in the second row correspond to the respective singular value spectra, indicating that UDV consistently seeks low-rank solutions regardless of the learning rate.
@@ B.4 Incorporating ReLU into UDV
We conducted a preliminary experiment to verify the feasibility of incorporating ReLU into the UDV, leaving a comprehensive study to future work. We derived the optimizers and learning rates from Tables SM3, SM4, SM5 and SM6, but only used the HPART dataset for regression and the MNIST for classification.
Figure SM23 and Figure SM24 showed that UDV-ReLU and UDV-ReLU-s exhibit a similar (or even more pronounced) decaying pattern in singular values as observed in the proposed UDV. The ‘stair-step’ accuracy curve was also observed in these models.
Figure SM23: ReLU in UDV: Singular value pattern on HPART dataset.
Figure SM24: ReLU in UDV: test accuracy (the first row) and the singular value pattern (the second row) on MNIST dataset. Faster convergence can be achieved by fine-tuning the learning rate.
@@ B.5 Comparison of UDV and Three-Layer Fully Connected Networks
The results, also included in Tables SM3, SM4, SM5 and SM6, are shown in Figure SM25 and Figure SM26. These findings demonstrate that the singular value decay in UDV unconstrained, UFV, and UV is significantly slower than in the proposed UDV framework. This underscores the critical role of the diagonal layer and constraints in facilitating low-rank solution identification within the proposed structure.
Figure SM25: Comparison of singular value spectrum among UDV, UDV (unconstrained),UFV and UV on the regression tasks.
Figure SM26: Comparison of singular value spectrum among UDV, UDV (unconstrained), UFV and UV on the classification datasets (MNIST).
@@ B.6 Comparing the Effects of UDV and Weight Decay Regularization
Increasing $\gamma$ led to a faster decay in the singular value spectrum but at the cost of higher training and test losses. Figure SM27 illustrates an example where we selected the largest $\gamma$ values for which the test loss remained within 10% of the UDV baseline. The resulting singular value decay in the UV and UFV models was less pronounced than in the UDV model. Further increasing $\gamma$ to match or exceed the spectral decay of the UDV model resulted in significantly worse test loss compared to the UDV network. The complete results are presented in Figures SM28 and SM29.
@@
Figure SM27: Comparison of singular value spectrum among the proposed UDV, UFV and UV (with weight decay) on the HPART dataset. In addition to the UDV baseline, the UFV and UV models include the regularization parameter $\gamma$ and the relative change in test performance. The relative change is defined as ($\frac{loss_{model}-loss_{baseline}}{loss_{baseline}}\times 100\%$), representing how much worse the model performs compared to the UDV baseline.
@@
Figure SM28: Extension of Figure SM27. Comparison of singular value spectrum between the proposed UDV and UFV (with weight decay) on the HPART dataset. Regularization parameter $\gamma=0$ indicates that no weight decay is applied.
@@
Figure SM29: Extension of Figure SM27. Comparison of singular value spectrum between the proposed UDV and UV (with weight decay) on the HPART dataset. Regularization parameter $\gamma=0$ indicates that no weight decay is applied.
B.7 A Toy Example of Causal Language Modeling Using a Pre-trained LLaMA-2 Model with LoRA-Based UDV Fine-Tuning
To further broaden the application scope of the proposed method, we integrate UDV into the pre-trained model using the LoRA framework and evaluate its effectiveness on the causal language modeling task (i.e., a generative task where a model predicts the next token in a sequence based solely on past tokens). Experiments are conducted using the LLaMA-2-7B model (Touvron et al., 2023) and the WikiText-2 dataset (Merity et al., 2016), which is a common benchmark test.
LoRA (Low-Rank Adaptation) is a parameter-efficient fine-tuning technique that inserts trainable low-rank matrices into pre-trained models (Hu et al., 2022). By freezing the original model weights and only training a small number of additional parameters, LoRA enables efficient adaptation to downstream tasks while significantly reducing computational and storage costs. Specifically, LoRA introduces trainable low-rank matrices $B\in\mathbb{R}^{d\times r}$ and $A\in\mathbb{R}^{r\times k}$ as a side path to the frozen pre-trained weight matrix $W_{0}$, and expresses the forward pass as $h=W_{0}x+BAx$, where $x$ is the input. Here we consider $d=k$ for simplicity.
By inserting a diagonal matrix $D$ between the matrices $B$ and $A$, the LoRA structure naturally extends to the LoRA-based UDV structure, as illustrated in Figure SM30. After the weight update of $U,D$ and $V$, constraints described in Equation 5 are applied to the LoRA-based UDV structure. In practice, the diagonal matrix can be efficiently implemented using a vector $d\in\mathbb{R}^{1\times r}$, further reducing computational and memory overhead. Consequently, the LoRA-based UDV introduces a negligible increase in the number of trainable parameters compared to the original LoRA structure, as it adds only $r$ parameters per LoRA layer replacement.
Figure SM30: LoRA forward pass (Left) and LoRA-based UDV forward pass (Right). The inputs and pre-trained weights are identical in both cases. Matrices with the same fill color indicate matching shapes.
We initialized both the LoRA and LoRA-based UDV structures from the same starting point: $B=V=0$, $A=U=\mathcal{N}(0,0.01)$ and $D=I$, where $I$ is the identity matrix. Mini-batch gradient descent (MBGD) with learning rates $lr=0.001$, using an effective batch size of 32, was applied to better observe the effects of the imposed constraints. Following the standard LoRA setup, only the linear layers associated with the Query ($Q$) and Value ($V$) projections in the attention mechanism were replaced with LoRA and LoRA-based UDV structures. We evaluated the models using rank values $r\in\{4,8,16,32,64\}$ , where the dimensionality of $d$ depends on the shape of the original layer. Both structures were fine-tuned for 30 epochs, and then pruning was applied as described in Section 3.1.3.
Instead of applying uniform pruning across all LoRA-based UDV structures, we set several thresholds on the retained energy—defined as the sum of the squared singular values—to dynamically prune each UDV block based on its individual importance. Note that the pruning can be applied to the LoRA structure in the same manner as the LoRA-based UDV structure by temporarily inserting an frozen identity matrix as a diagonal layer.
Perplexity (i.e., the exponential of the average negative log-likelihood) is used as the primary evaluation metric for this experiment, and lower perplexity indicates better language modeling performance. Furthermore, the pruning performance was demonstrated by the perplexity change in terms of the compression ratio (i.e., the ratio between the number of trainable parameters of the compressed structure and the original structure)
Since the hard constraints slower the convergence, and its fine-tuning performance is slightly lower than that of the original LoRA structure within the same number of epochs, as shown in Figure SM31 shown. However, the pruning results (see Figure SM32) demonstrate that the LoRA-based UDV is significantly more robust to high compression ratios. Even at a compression ratio of $0.1$, the LoRA-based UDV is able to maintain performance effectively.
Figure SM31: Comparison of perplexity between LoRA and LoRA-based UDV during training. The perplexity at epoch 0 reflects the performance of the pre-trained model before any fine-tuning.
Figure SM32: Comparison of pruning performance between LoRA and LoRA-based UDV during training.
This toy example is not intended to present a mature pruning method, but rather to highlight the potential of the proposed UDV block, which can be applied to a wide range of models and demonstrates promising pruning performance.
Appendix C Extended Discussion on the Related Work
Burer-Monteiro factorization. BM factorization was proposed for solving semidefinite programs (Burer & Monteiro, 2003, 2005; Boumal et al., 2016, 2020; Cifuentes, 2021) and has been recognized for its efficiency in addressing low-rank optimization problems, e.g., (Sun & Luo, 2016; Bhojanapalli et al., 2016; Park et al., 2017, 2018; Hsieh et al., 2018; Sahin et al., 2019; Lee et al., 2022; Yalçın et al., 2023). Building on the connections between training problems for (non-linear) two-layer neural networks and convex (copositive) formulations (see (Pilanci & Ergen, 2020; Ergen & Pilanci, 2020; Sahiner et al., 2021) and the references therein), Sahiner et al. (2023) recently introduced BM factorization to solve convex formulations of various architectures, including fully connected networks with ReLU and gated ReLU (Fiat et al., 2019) activations, two-layer convnets (LeCun et al., 1989a), and self-attention mechanisms based on Transformers (Vaswani et al., 2017).
Implicit regularization. One promising line of research that aims to explain the successful generalization abilities of neural networks follows the concept of ‘implicit regularization,’ which results from the optimization methods and formulations used during training (Neyshabur et al., 2014, 2017; Neyshabur, 2017). Several studies explore matrix factorization to investigate implicit bias (Gunasekar et al., 2017; Arora et al., 2018; Razin & Cohen, 2020; Belabbas, 2020; Li et al., 2021). Le & Jegelka (2022) extend these results to the final linear layers of nonlinear ReLU-activated feedforward networks with fully connected layers and skip connections. More recently Timor et al. (2023) investigated implicit regularization in ReLU networks. Much of the existing work focuses on gradient flow dynamics in the limit of infinitesimal learning rates. In particular, Gidel et al. (2019) examined discrete gradient dynamics in two-layer linear neural networks, showing that the dynamics progressively learn solutions of reduced-rank regression with a gradually increasing rank.
Constrained neural networks. Regularizers are frequently used in neural network training to prevent overfitting and improve generalization (Goodfellow et al., 2016), or to achieve structural benefits such as sparse and compact network architectures (Scardapane et al., 2017). However, it is conventional to apply these regularizers as penalty functions in the objective rather than as hard constraints, addressing them within gradient-based optimization via their (sub)gradients. This approach is likely favored due to the ease of implementation, as pre-built functions are readily available in common neural network packages. Several independent studies have also applied proximal methods across different networks and applications, finding that proximal-based methods tend to yield solutions with more pronounced structures (Bai et al., 2019; Yang et al., 2020; Yun et al., 2021; Yurtsever et al., 2021; Yang et al., 2022). Structural regularization in the form of hard constraints, however, appears to be rare in neural network training. One notable exception is in the context of neural network training with the Frank-Wolfe algorithm (Pokutta et al., 2020; Zimmer et al., 2022; Macdonald et al., 2022). Recently, Pethick et al. (2025) revealed parallels between Frank-Wolfe on constrained networks and algorithms that post-process update steps, such as Muon (Jordan et al., 2024), which achieves state-of-the-art results on nanoGPT by orthogonalizing the update directions before applying them.
There are other examples of constraints in neural networks. For instance, constraints can be applied to ensure adherence to real-world conditions in applications where they are necessary (Pathak et al., 2015; Márquez-Neila et al., 2017; Jia et al., 2017; Kervadec et al., 2019; Weber et al., 2021), or to incorporate physical laws in physics-informed neural networks (Raissi et al., 2019; Lu et al., 2021; Patel et al., 2022). In these cases, the feasible set is typically complex and difficult to project onto; therefore, optimization algorithms like primal-dual methods or augmented Lagrangian techniques are used. Constraints are also present in lifted neural networks, a framework where the training problem is reformulated in a higher-dimensional ‘lifted’ space. In this space, conventional activation functions like ReLU, used in the original formulation, are expressed as constraints (Askari et al., 2018; Sahiner et al., 2021; Bartan & Pilanci, 2021; Prakhya et al., 2024).
Pruning. Neural networks are overparameterized, which can enhance generalization and avoid poor local minima. But such models then suffer from excessive memory and computational demands, making them less efficient for deployment in real-world applications (Chang et al., 2021). Several compression techniques have been studied in the literature, including parameter quantization (Krishnamoorthi, 2018), knowledge distillation (Gou et al., 2021), and pruning.
Pruning reduces the number of parameters, resulting in more compact and efficient models that are easier to deploy. The literature on pruning is extensive, with diverse methods proposed with distinct characteristics. Various criteria are used to determine which weights to prune, including second-order derivative-based methods (LeCun et al., 1989b; Hassibi & Stork, 1992), magnitude-based pruning (Janowsky, 1989; Han et al., 2015), saliency heuristics (Mozer & Smolensky, 1988; Lee et al., 2018), and matrix or tensor factorization-based techniques (Xue et al., 2013; Sainath et al., 2013; Jaderberg et al., 2014; Lebedev et al., 2015; Swaminathan et al., 2020), among others. Pruning by singular value thresholding has shown promising results, particularly in natural language processing (Chen et al., 2021), and is often used along with various enhancements such as importance weights and data whitening for effective compression of large language models (Hsu et al., 2022; Yuan et al., 2023; Wang et al., 2024). A comprehensive review of pruning methods is beyond the scope of this paper due to space limitations and the diversity of approaches. For more detailed reviews, we refer to (Reed, 1993; Blalock et al., 2020; Cheng et al., 2024), and the references therein.
@@ Appendix C Extended Discussion on the Related Work
Table SM3: Experiments using Adam optimizer. Not applicable or results with obvious oscillations or divergence are denoted as ‘–’.
@@
(MaxVit-T [M] $|$ EfficientNet-B0 [E] $|$ RegNetX-32GF [R])
($\times 100\%$)
@@
Table SM4: Experiments using NAdam optimizer. Not applicable or results with obvious oscillations or divergence are denoted as ‘–’.
@@
(MaxVit-T [M] $|$ EfficientNet-B0 [E] $|$ RegNetX-32GF [R])
($\times 100\%$)
@@
Table SM5: Experiments using MBGD optimizer. Not applicable or results with obvious oscillations or divergence are denoted as ‘–’.
@@
(MaxVit-T [M] $|$ EfficientNet-B0 [E] $|$ RegNetX-32GF [R])
($\times 100\%$)
@@
Table SM6: Experiments using MBGDM optimizer. Not applicable or results with obvious oscillations or divergence are denoted as ‘–’.
@@
(MaxVit-T [M] $|$ EfficientNet-B0 [E] $|$ RegNetX-32GF [R])
($\times 100\%$)